OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). (sin x)(cos x) = 0 A. pi/2, pi B. 0, pi/2, pi, 3pi/2 C. pi, 3pi/2 D. 0, 3pi/2

OpenStudy (solomonzelman):

`cos(x) = 0` when? `sin(x) = 0` when?

OpenStudy (solomonzelman):

cos(0) is not zero, but sin(0) is zero. So your first solution is x=0. can you tell me some more solutions?

OpenStudy (anonymous):

Given the options, 3pi/2 is a solution. However I'm not exactly sure how.. I'm lost on this one.

OpenStudy (solomonzelman):

\(\large\color{black}{ 3\pi/2 }\) is the same as \(\large\color{black}{ 270 }\) degrees. and what is zero, sin(270) or cos(270) ?

OpenStudy (anonymous):

sin(270)= -1 cos(270)= 0

OpenStudy (anonymous):

so cos(270)

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

This is why \(\large\color{black}{ 3\pi/2 }\) is one of the solutions.

OpenStudy (solomonzelman):

okay, and knowing (from what we said before) that zero is one of the solutions, you now have \(\large\color{black}{ B }\) or \(\large\color{black}{ D }\), right?

OpenStudy (solomonzelman):

Then, try \(\large\color{black}{ \pi/2 }\) and \(\large\color{black}{ \pi ~. }\)

OpenStudy (solomonzelman):

1) the sine or the cosine of \(\large\color{black}{ \pi/2 }\) is zero? 2) the sine or the cosine of \(\large\color{black}{ \pi }\) is zero?

OpenStudy (anonymous):

sin(pi/2)= 1 cos(pi/2)= 0 ** sin(pi)= 0 ** cos(pi)= -1

OpenStudy (anonymous):

1) cosine 2) sine

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

And which option will you pick as your final answer?

OpenStudy (anonymous):

B?

OpenStudy (solomonzelman):

Yup:)

OpenStudy (anonymous):

Thank you so much! :)

OpenStudy (solomonzelman):

No problem.